\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{schnitzer2019clifford}
H.J.~Schnitzer, \emph{Clifford group and stabilizer states from
  {C}hern-{S}imons theory},
  [\href{https://arxiv.org/abs/arXiv:1903.06789}{{\ttfamily
  arXiv:1903.06789}}].

\bibitem{schnitzer2020n}
H.J.~Schnitzer, \emph{{$\mathrm{SU}(N)_1$} {C}hern-{S}imons theory, the
  {C}lifford group, and entropy cone},
  [\href{https://arxiv.org/abs/arXiv:2008.02406}{{\ttfamily
  arXiv:2008.02406}}].

\bibitem{schnitzer2020clifford}
H.J.~Schnitzer, \emph{Clifford operators in {$\mathrm{SU}(N)_1$}; {$N$} not odd
  prime},  [\href{https://arxiv.org/abs/arXiv:2011.06035}{{\ttfamily
  arXiv:2011.06035}}].

\bibitem{steinhoff2017qudit}
F.E.~Steinhoff, C.~Ritz, N.~Miklin and O.~G{\"u}hne, \emph{Qudit hypergraph
  states}, {\emph{Physical Review A} {\bfseries 95} (2017) 052340}
  [\href{https://arxiv.org/abs/arXiv:1612.06418}{{\ttfamily
  arXiv:1612.06418}}].

\bibitem{englbrecht2020symmetries}
M.~Englbrecht and B.~Kraus, \emph{Symmetries and entanglement of stabilizer
  states}, {\emph{Physical Review A} {\bfseries 101} (2020) 062302}
  [\href{https://arxiv.org/abs/arXiv:2001.07106}{{\ttfamily
  arXiv:2001.07106}}].

\bibitem{liu2020many}
Z.-W.~Liu and A.~Winter, \emph{Many-body quantum magic},
  [\href{https://arxiv.org/abs/arXiv:2010.13817}{{\ttfamily
  arXiv:2010.13817}}].

\bibitem{rossi2013quantum}
M.~Rossi, M.~Huber, D.~Bru{\ss} and C.~Macchiavello, \emph{Quantum hypergraph
  states}, {\emph{New Journal of Physics} {\bfseries 15} (2013) 113022}
  [\href{https://arxiv.org/abs/arXiv:1211.5554}{{\ttfamily arXiv:1211.5554}}].

\bibitem{hein2006entanglement}
M.~Hein, W.~D{\"u}r, J.~Eisert, R.~Raussendorf, M.~Nest and H.-J.~Briegel,
  \emph{Entanglement in graph states and its applications},
  [\href{https://arxiv.org/abs/quant-ph/0602096}{{\ttfamily
  quant-ph/0602096}}].

\bibitem{mlawer1991group}
E.J.~Mlawer, S.G.~Naculich, H.A.~Riggs and H.J.~Schnitzer, \emph{Group-level
  duality of {WZW} fusion coefficients and {C}hern-{S}imons link observables},
  {\emph{Nuclear Physics B} {\bfseries 352} (1991) 863}.

\bibitem{salton2017entanglement}
G.~Salton, B.~Swingle and M.~Walter, \emph{Entanglement from topology in
  {C}hern-{S}imons theory}, {\emph{Physical Review D} {\bfseries 95} (2017)
  105007} [\href{https://arxiv.org/abs/arXiv:1611.01516}{{\ttfamily
  arXiv:1611.01516}}].

\bibitem{heyfron2019quantum}
L.E.~Heyfron and E.~Campbell, \emph{A quantum compiler for qudits of prime
  dimension greater than 3},
  [\href{https://arxiv.org/abs/arXiv:1902.05634}{{\ttfamily
  arXiv:1902.05634}}].

\bibitem{paetznick2013universal}
A.~Paetznick and B.W.~Reichardt, \emph{Universal fault-tolerant quantum
  computation with only transversal gates and error correction},
  {\emph{Physical review letters} {\bfseries 111} (2013) 090505}
  [\href{https://arxiv.org/abs/arXiv:1304.3709}{{\ttfamily arXiv:1304.3709}}].

\bibitem{o2017quantum}
J.~O'Gorman and E.T.~Campbell, \emph{Quantum computation with realistic
  magic-state factories}, {\emph{Physical Review A} {\bfseries 95} (2017)
  032338} [\href{https://arxiv.org/abs/arXiv:1605.07197}{{\ttfamily
  arXiv:1605.07197}}].

\bibitem{eastin2013distilling}
B.~Eastin, \emph{Distilling one-qubit magic states into {T}offoli states},
  {\emph{Physical Review A} {\bfseries 87} (2013) 032321}
  [\href{https://arxiv.org/abs/arXiv:1212.4872}{{\ttfamily arXiv:1212.4872}}].

\bibitem{jones2013low}
C.~Jones, \emph{Low-overhead constructions for the fault-tolerant {T}offoli
  gate}, {\emph{Physical Review A} {\bfseries 87} (2013) 022328}
  [\href{https://arxiv.org/abs/arXiv:1212.5069}{{\ttfamily arXiv:1212.5069}}].

\bibitem{howard2012qudit}
M.~Howard and J.~Vala, \emph{Qudit versions of the qubit $\pi$/8 gate},
  {\emph{Physical Review A} {\bfseries 86} (2012) 022316}
  [\href{https://arxiv.org/abs/arXiv:1206.1598}{{\ttfamily arXiv:1206.1598}}].

\bibitem{biswal2019techniques}
L.~Biswal, D.~Bhattacharjee, A.~Chattopadhyay and H.~Rahaman, \emph{Techniques
  for fault-tolerant decomposition of a multicontrolled {T}offoli gate},
  {\emph{Physical Review A} {\bfseries 100} (2019) 062326}
  [\href{https://arxiv.org/abs/arXiv:1904.06920}{{\ttfamily
  arXiv:1904.06920}}].

\bibitem{haah2018codes}
J.~Haah and M.B.~Hastings, \emph{Codes and protocols for distilling {$t$},
  controlled-{$s$}, and {T}offoli gates}, {\emph{Quantum} {\bfseries 2} (2018)
  71} [\href{https://arxiv.org/abs/arXiv:1709.02832}{{\ttfamily
  arXiv:1709.02832}}].

\bibitem{fliss2020knots}
J.R.~Fliss, \emph{Knots, links, and long-range magic},
  [\href{https://arxiv.org/abs/arXiv:2011.01962}{{\ttfamily
  arXiv:2011.01962}}].

\bibitem{balasubramanian2017multi}
V.~Balasubramanian, J.R.~Fliss, R.G.~Leigh and O.~Parrikar,
  \emph{Multi-boundary entanglement in {C}hern-{S}imons theory and link
  invariants}, {\emph{Journal of High Energy Physics} {\bfseries 2017} (2017)
  61} [\href{https://arxiv.org/abs/arXiv:1611.05460}{{\ttfamily
  arXiv:1611.05460}}].

\bibitem{balasubramanian2018entanglement}
V.~Balasubramanian, M.~DeCross, J.~Fliss, A.~Kar, R.G.~Leigh and O.~Parrikar,
  \emph{Entanglement entropy and the colored {J}ones polynomial},
  {\emph{Journal of High Energy Physics} {\bfseries 2018} (2018) 1}
  [\href{https://arxiv.org/abs/arXiv:1801.01131}{{\ttfamily
  arXiv:1801.01131}}].

\bibitem{Jain_2020}
A.~Jain and S.~Prakash, \emph{Qutrit and ququint magic states},
  \href{https://doi.org/10.1103/physreva.102.042409}{\emph{Physical Review A}
  {\bfseries 102} (2020) }.

\bibitem{white2020conformal}
C.D.~White, C.~Cao and B.~Swingle, \emph{Conformal field theories are magical},
   [\href{https://arxiv.org/abs/arXiv:2007.01303}{{\ttfamily
  arXiv:2007.01303}}].

\bibitem{white2020mana}
C.D.~White and J.H.~Wilson, \emph{Mana in {H}aar-random states},
  [\href{https://arxiv.org/abs/arXiv:2011.13937}{{\ttfamily
  arXiv:2011.13937}}].

\bibitem{veitch2014resource}
V.~Veitch, S.H.~Mousavian, D.~Gottesman and J.~Emerson, \emph{The resource
  theory of stabilizer quantum computation}, {\emph{New Journal of Physics}
  {\bfseries 16} (2014) 013009}
  [\href{https://arxiv.org/abs/arXiv:1307.7171}{{\ttfamily arXiv:1307.7171}}].

\bibitem{freedman2002modular}
M.H.~Freedman, M.~Larsen and Z.~Wang, \emph{A modular functor which is
  universal for quantum computation}, {\emph{Communications in Mathematical
  Physics} {\bfseries 227} (2002) 605}
  [\href{https://arxiv.org/abs/quant-ph/0001108}{{\ttfamily
  quant-ph/0001108}}].

\bibitem{schnitzer2018level}
H.J.~Schnitzer, \emph{Level-rank duality for a universal topological quantum
  computer},  [\href{https://arxiv.org/abs/arXiv:1811.11861}{{\ttfamily
  arXiv:1811.11861}}].

\bibitem{bao2020quantum}
N.~Bao, N.~Cheng, S.~Hern{\'a}ndez-Cuenca and V.P.~Su, \emph{The quantum
  entropy cone of hypergraphs},
  [\href{https://arxiv.org/abs/arXiv:2002.05317}{{\ttfamily
  arXiv:2002.05317}}].

\bibitem{walter2020hypergraph}
M.~Walter and F.~Witteveen, \emph{Hypergraph min-cuts from quantum entropies},
  [\href{https://arxiv.org/abs/arXiv:2002.12397}{{\ttfamily
  arXiv:2002.12397}}].

\bibitem{bao2020gap}
N.~Bao, N.~Cheng, S.~Hern{\'a}ndez-Cuenca and V.P.~Su, \emph{A gap between the
  hypergraph and stabilizer entropy cones},
  [\href{https://arxiv.org/abs/arXiv:2006.16292}{{\ttfamily
  arXiv:2006.16292}}].

\bibitem{bao2020bit}
N.~Bao and J.~Harper, \emph{Bit threads on hypergraphs},
  [\href{https://arxiv.org/abs/arXiv:2012.07872}{{\ttfamily
  arXiv:2012.07872}}].

\end{thebibliography}\endgroup

\documentclass[12pt]{article}
\pdfoutput=1
\usepackage{euscript,amsmath,amsthm,amsfonts,amssymb}
\usepackage{mathtools}
\usepackage{subfig}
\usepackage{float}
\usepackage[margin=1in]{geometry}
\usepackage{graphicx}
\usepackage{outlines}
\usepackage[dvipsnames]{xcolor}
\usepackage{jheppub}
\usepackage{cleveref}

\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{todonotes}
\usepackage{listings}

\usepackage[utf8]{inputenc}
\usepackage{color}
\usepackage{graphicx}
\usepackage{caption}
\usepackage{empheq}

\newcommand{\bra}[1]{\left\langle#1\right|}
\newcommand{\ket}[1]{\left|#1\right\rangle}

\newcommand{\su}[1]{\mathrm{SU}(#1)}
\newcommand{\sun}{\su{N}_1}

\newcommand{\nmod}{5 \mod{4}}
\newcommand{\mvec}[1]{\mathbf{#1}}

\numberwithin{equation}{section}

\title{Hypergraph States in $\su{N}_1$, $N$ odd prime, Chern-Simons Theory}
\author{Howard J.~Schnitzer}
\affiliation{Martin Fisher School of Physics, Brandeis University, Waltham MA 02453}
\preprint{BRX-TH-6672}
\emailAdd{schnitzr@brandeis.edu}

\abstract{Entropy cones for $SU(N)_{1}$ Chern-Simons theory are discussed. It is shown that stabilizer states can be constructed from topological operators in $SU(N)_{1}$ for $N$ odd prime, but not for $SU(N)_{K}$; $K \geq 2$. This implies that the topological entropy cone is properly contained in the stabilizer entropy cone for $SU(N)_{K}$; $K \geq 2$.

}

\abstract{Graph states and hypergraph states can be constructed from products of basic operations that appear in $\su{N}_1$. The level-rank dual of a theorem of Salton, Swingle, and Walter implies that these operations can be prepared topologically in the $n$-torus Hilbert space of Chern-Simons theory for $N \neq 5 \mod{4}$. 
    
For $\su{N}_1$, $N = 5 \mod{4}$, only stabilizer states can be prepared on the $n$-torus Hilbert space, which restricts the construction to graph states.}

\begin{document}
\maketitle
\flushbottom

\section{Introduction}
\label{sec:sec1}

The generalized Pauli group and Clifford operators \cite{schnitzer2019clifford,schnitzer2020n,schnitzer2020clifford} are obtained from products of basic operations of $\sun$ Chern-Simons theory for $N$ odd prime. Graph and hypergraph states \cite{steinhoff2017qudit,englbrecht2020symmetries,liu2020many,rossi2013quantum,hein2006entanglement} are described as products of such operations, which allow for explicit representations of graph and hypergraph states. For $N$ odd prime, $N \neq \nmod$, these operations can be obtained topologically in the $n$-torus Hilbert space of Chern-Simons theory as a result of the level rank dual \cite{mlawer1991group} of a theorem of Salton, Swingle, and Walter \cite{salton2017entanglement}. Thus graph and hypergraph states of $\sun$, $N \neq \nmod$, are topological on the $n$-torus Hilbert space of Chern-Simons theory.

For $\sun$, $N = \nmod$, only graph states are topological, as in this case only stabilizer states are obtained from the $n$-torus Hilbert space. Hypergraph states for $\sun$, $N = \nmod$, are therefore not topological in the above sense.

In Section \ref{sec:sec2} the construction of the generalized Pauli group and Clifford operations for $\sun$ is reviewed. Section \ref{sec:sec3} presents the construction of graph and hypergraph states in terms of basic operations of $\sun$. Section \ref{sec:sec4} is a discussion of related issues.

\section{$\su{d}_1$, $d$ odd prime}
\label{sec:sec2}

We first review the generalized Pauli group and Clifford operations for $\su{d}_1$, $d$ odd prime, following \cite{schnitzer2019clifford,schnitzer2020n,schnitzer2020clifford}.

\subsection{The $\su{d}_1$ Pauli group}
\label{sec:sec2a}

Representations of $\su{d}_1$ describing qudits are given by a single column Young tableau, with zero, one, \ldots, $(d-1)$ boxes. The fusion tensor of the theory is
\begin{equation}
    \label{eq:fusion}
    N_{ab}{}^c; \quad \quad a+b = c \mod{d}
\end{equation}
so that
\begin{equation}
    \label{eq:fusion2}
    N\ket{a}\ket{b} = \ket{a} \ket{a+b, \mod{d}}.
\end{equation}
The modular transformation matrix $S_{ab}$ satisfies
\begin{equation}
    \ket{a} = \sum_{b = 0}^{d-1} S_{ab} \ket{b}, \quad \quad a = 0 \ldots d-1.
\end{equation}
For $\omega$ a primitive $d$th root of unity
\begin{equation}
    \omega = \exp \left( \frac{2 \pi i}{d} \right)
\end{equation}
so that
\begin{equation}
    \label{eq:qft}
    S^* = \frac{1}{\sqrt{d}} \sum_{a = 0}^{d-1} \sum_{b = 0}^{d-1} \omega^{ab} \ket{a} \bra{b}
\end{equation}
which is the $d$-dimensional generalization of the Hadamard gate.

The Pauli operator $Z$ is given by
\begin{equation}
    Z_{ac} = \sum_{b = 0}^{d-1} S_{ab} \left( S^{\dagger}_{b+1,a} \right) \delta_{ac}
\end{equation}
so that
\begin{equation}
    Z = \sum_{a = 0}^{d-1} \omega^a \ket{a} \bra{a}.
\end{equation}
The Pauli operator $X$ is obtained from the fusion matrix, since
\begin{equation}
    \label{eq:paulix}
    N_{a,1}{}^b\ket{a} = \ket{a+1, \mod{d}}
\end{equation}
which is identical to
\begin{equation}
    X \ket{a} = \ket{a+1, \mod{d}}
\end{equation}
The single qudit Pauli group is the collection of operators
\begin{equation}
    \omega^r X^a Z^b; \quad \quad a,b,r \in \mathbb{Z}_d.
\end{equation}
Thus the one-qudit Pauli group is constructed from basic operations of $\su{d}_1$, $d$ odd.

The $n$-qudit Pauli group is obtained from products of operators of the one-qudit Pauli group. That is
\begin{equation}
    X^{\mvec{a}} Z^{\mvec{b}} = X^{a_1} Z^{b_1} \otimes X^{a_2} Z^{b_2} \otimes \cdots \otimes X^{a_n} Z^{b_n}
\end{equation}
The operator $X^{\mvec{a}} Z^{\mvec{b}}$, along with all scalar multiples thereof,
\begin{equation}
    \left\{ \omega^c X^{\mvec{a}} Z^{\mvec{b}} \ | \ c \in \mathbb{Z}_d \right\}
\end{equation}
defines the $n$-qudit Pauli group.

\subsection{$\su{d}_1$ Clifford operators, $d$ odd}

The necessary gates for the single-qudit Clifford operators are \cite{schnitzer2019clifford,schnitzer2020n,schnitzer2020clifford} i) the QFT gate, Eq.~\eqref{eq:qft}, and ii) the phase gate
\begin{equation}
    P \ket{j} = \omega^{j(j-1)/2} \ket{j}.
\end{equation}
The multi-qudit Clifford operators are obtained from the generalizations of \eqref{eq:qft} and \eqref{eq:paulix} as well as the SUM gate,
\begin{align}
    C_{\mathrm{SUM}} \ket{i} \ket{j} &= N \ket{i} \ket{j} \nonumber \\
    &= \ket{i} \ket{i+j, \mod{d}}
\end{align}

\section{Graph and hypergraph states}
\label{sec:sec3}

\subsection{Graph states}
\label{sec:sec3a}

There are many equivalent constructions of graph states \cite{steinhoff2017qudit,englbrecht2020symmetries,liu2020many,rossi2013quantum,hein2006entanglement}. We follow arxiv:1612.06418 for a definition of qudit graph states. The multigraph is $G = (V,E)$, with vertices $V$ and edges $E$, where an edge has multiplicity $m_e \in \mathbb{Z}_d$. To $G$ associate a state $\ket{G}$ such that to each vertex $i \in V$, there is a local state
\begin{equation}
    \ket{+} = \ket{p_0} = \frac{1}{\sqrt{d}} \sum_{q = 0}^{d-1} \ket{q}
\end{equation}
Recall that the Hadamard gate generalizes to \eqref{eq:qft}, so that
\begin{align}
    S^* \ket{0} &= \frac{1}{\sqrt{d}} \sum_{q = 0}^{d-1} \ket{q} \nonumber \\
    &= \ket{+} = \ket{p_0}.
\end{align}

To each edge $e = \{i,j\}$ apply the unitary
\begin{equation}
    Z_e^{m_e} = \sum_{q_i = 0}^{d-1} \ket{q_i} \bra{q_i} \otimes \left(Z_j^{m_e}\right)^{q_i}
\end{equation}
to the state
\begin{equation}
    \ket{+}^V = \bigotimes_{i \in V} \ket{+}_i
\end{equation}

The graph state is
\begin{align}
    \ket{G} &= \prod_{e \in E} Z_e^{m_e} \ket{+}^V \\
        &= \prod_{e \in E} Z_e^{m_e} \bigotimes_{i \in V} \ket{+}_i
\end{align}
The level-rank dual \cite{mlawer1991group} of Theorem 1 of Salton, Swingle, and Walter \cite{salton2017entanglement} for $d$ odd prime implies that the graph state $\ket{G}$ can be constructed from topological operations on the $n$-torus Hilbert space of Chern-Simons $\su{d}_1$ by means of the operations detailed in Section \ref{sec:sec2}. Every stabilizer state is LC equivalent to a graph state, while the Clifford group enables conversion between different multigraphs \cite{steinhoff2017qudit,englbrecht2020symmetries,liu2020many,rossi2013quantum,hein2006entanglement}.

\subsection{Hypergraph states}

We again follow arxiv:1612.06418 for the construction of qudit multi-hypergraph states. Given a multi-hypergraph $H = (V,E)$, associate a quantum state $\ket{H}$, with $m_e \in \mathbb{Z}_d$ the multiplicity of the hyperedge $e$. To each vertex $i \in V$, associate a local state
\begin{align}
    \ket{+} &= \frac{1}{\sqrt{d}} \sum_{q=0}^{d-1} \ket{q} \nonumber \\
        &= S^* \ket{0}
\end{align}
To each hyperedge $e \in E$, with multiplicity $m_e$, apply the controlled unitary $Z_e^{m_e}$ to the state
\begin{equation}
    \ket{+}^V = \bigotimes_{i \in V} \ket{+}_i
\end{equation}
The hypergraph state is
\begin{align}
    \ket{H} = \prod_{e \in E} Z_e^{m_e} \ket{+}^V
\end{align}
The elementary hypergraph state is
\begin{align}
    \ket{H} = \sum_{q=0}^{d-1} \ket{q} \bra{q} \otimes \left( Z_{e \backslash \{1\}}^{m_e} \right)^q \ket{+}^V
\end{align}
For $d$ prime, all $n$-elementary hypergraph states are equivalent under SLOCC.

Hypergraph and graph states admit a representation in terms of Boolean functions,
\begin{equation}
    \ket{H} = \sum_{q = 0}^{d-1} \omega^{f(q)} \ket{q}
\end{equation}
with $f: \mathbb{Z}_d^n \rightarrow \mathbb{Z}_d$, where
\begin{equation}
    \label{eq:boolfunc}
    f(x) = \sum_{\substack{i_1,\ldots,i_k \in V \\ \{ i_1,\ldots,i_k \} \in E}} x_{i_1} \cdots x_{i_k}
\end{equation}
For graph states, $f(x)$ is quadratic, i.e.
\begin{equation}
    f(x) = \sum_{\substack{i_1,i_2 \in V \\ \{ i_1,i_2 \} \in E}} x_{i_1} x_{i_2}
\end{equation}
while for $f(x)$ cubic or higher, $\ket{H}$ is a hypergraph state. Therefore, for quadratic $f(x)$, one has a representation of stabilizer states, up to LC equivalence. For $f(x)$ cubic or higher, $\ket{H}$ represents hypergraph states which contain ``magic'' states. Examples of magic states are the $\mathrm{CCZ}$ state and Toffoli states, constructed from appropriate gates \cite{heyfron2019quantum,paetznick2013universal,o2017quantum,schnitzer2020clifford,eastin2013distilling,jones2013low,howard2012qudit,biswal2019techniques,haah2018codes}. Thus
\begin{equation}
    \mathrm{CCZ} \ket{x_1 x_2 x_3} = \omega^{x_1 x_2 x_3} \ket{x_1 x_2 x_3}
\end{equation}
with
\begin{equation}
    \ket{\mathrm{CCZ}} = \mathrm{CCZ} \ket{+ \otimes^3}
\end{equation}
as an example of a magic hypergraph state. Similarly
\begin{equation}
    \ket{\mathrm{Toff}} = \mathrm{Toff} \ket{+ \otimes^3}
\end{equation}
where the Toffoli gate can be expressed in terms of the fusion matrix \eqref{eq:fusion2}. Explicitly,
\begin{align}
    \mathrm{Toff} \ket{i,j,k} &= N_{ij,k}{}^{(ij+k)} \nonumber \\
        &= \ket{i,j,ij+k, \mod{d}}
\end{align}
where the Young tableau for $(ij)$ has $i+j$ vertical boxes, $\mod{d}$.

For $\su{d}_1$, $d$ odd prime, $d \neq \nmod$, both graph and hypergraph states can be obtained from operators which can be constructed from products of topological operations on the $n$-torus Hilbert space \cite{salton2017entanglement}.

For $\su{d}_1$, $d = \nmod$, the topological argument does not apply, since in this case Theorem 1 of Salton \emph{et al} \cite{salton2017entanglement} implies that \emph{only} stabilizer states can be constructed on the $n$-torus Hilbert space. Thus, graph states can be so constructed but not hypergraph states with cubic or higher functions \eqref{eq:boolfunc}.

\section{Discussion}
\label{sec:sec4}

It was shown above that graph states and hypergraph states for $\su{d}_1$, $d$ odd, can be constructed from the basic operations $N_{ab}{}^c$, $S_{ab}$, $Z$, and $X$ of $\su{d}_1$ Chern-Simons theory. For $d$ odd prime, $d \neq \nmod$, these operations can be constructed topologically on the $n$-torus Hilbert space of Chern-Simons theory. A subset of hypergraph states are ``magic.'' For $d \neq \nmod$, they are topological in the above sense.

Fliss \cite{fliss2020knots} has studied knot and link states of $\su{2}_d$ Chern-Simons theory, and has shown that knot and link states are generically magical. However for $\mathrm{U}(1)_d$, magic is absent for all knot and link states. Since $\mathrm{U}(1)_d$ is level-rank dual to $\su{d}_1$, the knot and link states for this theory also have zero magic \cite{fliss2020knots,balasubramanian2017multi,balasubramanian2018entanglement},\cite{Jain_2020}.

There is a great deal of recent interest in magic states \cite{white2020conformal,white2020mana,liu2020many,veitch2014resource}. One feature that deserves further study is to understand which magic states are topological. For example, universal topological computing is possible for $\su{2}_3$ \cite{freedman2002modular} and $\su{3}_2$ \cite{schnitzer2018level} Chern-Simons theory. Implicitly this implies that magic states are present in these theories, presumably due to the braiding operations. It would be interesting to make this explicit.

\acknowledgments
\addcontentsline{toc}{section}{Acknowledgments}

We thank Zi-Wen Liu for clarifying the difference between hypergraph / stabilizer cones \cite{bao2020quantum,walter2020hypergraph,bao2020gap,bao2020bit}, and hypergraph / stabilizer states, and for emphasizing that universal quantum computation models for $\su{2}_3$ and $\su{3}_2$ have gates / actions which are magical.

We are grateful to Greg Bentsen and Isaac Cohen for their assistance in the preparation of the paper.

\addcontentsline{toc}{section}{References}

\nocite{*}
\bibliographystyle{JHEP}
\bibliography{main}

\end{document}